SOLUTION OF CALCULATING THE EQUIVALENT WEIGHT OF MOHR’S SALT
 

Grade:11

1 Answers

Bhavya
7 years ago
Dear student
n-factor for a salt is either total cationic charge/total anionic charge excluding the charge on cation/anion of water.
Mohr’s salt(NH₄)₂Fe(SO₄)₂·6H₂O
Here total cationic charge (2 NH4+ and 1 Fe+2) = total anionic charge (2 SO4-2) = 4
Equivalent weight = Molar mass/ n-factor = 392/ 4 = 98